# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 3.6 Microscopic CDCC
The standard CDCC method assumes a cluster (two-body or three-body) description of the projectile nucleus. This simplification has of course limitations and drawbacks. For example: (i) it requires cluster-target optical potentials, which are are not always well determined; (ii) the extension to more than three bodies is very challenging and currently not available; (iii) excitations of the fragments are ignored altogether or, at most, approximately included with some collective model. To overcome these problems, a microscopic version of the CDCC method (MCDCC) has been proposed by Descouvemont and co-workers Des13 ; Des18 . The method uses a many-body description of the projectile states, based on a cluster approximation, known as resonating group method (RGM). In the RGM, an eigenstate of the projectile Hamiltonian is written as an antisymmetric product of cluster wave functions. For example, for a 7Li projectile, described as $\alpha+t$, the RGM wave function is expressed as:
@@
where $\phi_{\alpha}$ and $\phi_{\alpha}$ are shell model wave functions of the $\alpha$ and $t$ clusters, $\ell$ their relative orbital angular momentum and $J$ the total spin. In Eq. (28), $\vec{\rho}$ is the relative coordinate (see Fig. 15), and ${\mathcal{A}}$ is the 7-body antisymmetrization operator which takes into account the Pauli principle among the 7 nucleons of the projectile. The function $g^{\ell J}_{i}(\rho)$ is determined from a Schrödinger equation associated with the projectile Hamiltonian. Continuum states are included using a pseudo-state basis.
@@
Figure 15 shows an application of the method to the reaction 7Li+208Pb at near-barrier energies. Experimental data are compared with a one-channel calculation (only 7Li g.s.), two-channels calculations (ground plus first excited state) and several CDCC calculations including continuum states up to a certain projectile angular momentum. It is seen that a good description of the data is achieved when a sufficiently large number of continuum states is included.
@@
Other applications of the microscopic CDCC method include its extension for three-body projectiles including a single nucleon as one of the three particles, such as 8Li or 8B as $(^{4}\text{He}+t+n)$ and $(^{4}\text{He}+{{}^{3}}\text{He}+p)$, respectively Descouvemont2018 . Figure 16 shows results for the elastic scattering of 7Li on 12C and 209Bi targets at low energies. The dashed line corresponds to the calculation including only the projectile ground state, while the solid line are the full microscopic CDCC calculations that exhibit a better agreement with the data.
@@ 3.7 Interpretation of the CDCC wavefunction and cross sections
For a meaningful comparison with the experimental data, it is important to properly understand which information and observables are (and which are not!) provided by the CDCC wavefunction. Of course, elastic scattering is the most direct outcome of the method and the one that can be compared with the data more easily. By construction, the elastic scattering produced by the CDCC calculation incorporates the effect of the coupling to the inelastic channels and the breakup channels included in the CDCC modelspace. According to our former notation, these breakup channels correspond to the three-body final states of the form $b+x+A$ in which the target and the projectile subsystems $b$ and $x$ escape the interaction region and remain in their ground state. Morover, in addition to these channels explicitly included in the modelspace, the CDCC includes effectively the effects of other channels associated with possible excitations of the target and/or of the fragments via the optical potentials.
@@
When the binning method is employed, a double differential cross section (with respect to the c.m. angle and relative energy) can be obtained by dividing each angular distribution by the bin width, as follows:
@@
Regarding the nonelastic breakup contributions, which are effectively included in the fragment-target imaginary potentials, the CDCC method by itself does not provide detailed breakup cross sections involving excitations. However, one can define and compute a total absorption cross section that accounts for their overall contribution. For that, one first derives a reaction cross section in terms of the CDCC S-matrices asnve [^1] See Ref. Des17 for a different convention.:
$\sigma_{R}=\frac{\pi}{k^{2}_{i}}\sum_{J_{T}\pi}\sum_{L_{i}}\frac{2J_{T}+1}{(2J^{i}_{p}+1)(2J^{i}_{t}+1)}\left(1-|S^{J_{T}}_{\beta_{i},\beta_{i}}|^{2}\right),$
@@
$\sigma_{\text{bu}}=\frac{\pi}{k^{2}_{i}}\frac{k}{k_{i}}\sum_{J_{T}\pi}\sum_{LL_{i}}\frac{2J_{T}+1}{(2J^{i}_{p}+1)(2J^{i}_{t}+1)}|S^{J_{T}}_{\beta,\beta_{i}}|^{2}.$
Then, one may define an absorption cross section as $\sigma_{\text{abs}}=\sigma_{\text{R}}-\sigma_{\text{bu}}$. This absorption cross section will account for all nonelastic processes associated with the absorptive part of the fragment-target potentials, such as target excitation (i.e., inelastic scattering), complete fusion (i.e., capture of the whole projectile) or incomplete fusion (i.e., capture of one of the projectile fragments). Very often, comparison with experimental data requires more detailed cross sections, such as the angular or energy distributions of the fragment that survives after the capture (absorption) of the other fragment. As noted above, these observables are not directly provided by the CDCC method, at least in its standard formulation. However, as will be shown in Sect. 3.7.2, the CDCC wavefunction can be used as an input for a suitable formalism capable of providing such observables.
The absorption cross section plays an important role in understanding the CDCC results. To quantify the importance of the continuum on the elastic cross sections it is common to compare full CDCC results with no-continuum calculations, in which only the ground-state to ground-state coupling potential is retained. An example is shown in Fig. 17, which depicts the results for the 6He+208Pb reaction at 18 MeV and the 8B+90Zr reaction at 26.5 MeV. Dashed and solid lines denote the no-continuum and full CDCC calculations, respectively (in the 6He case, the improved dineutron model of Ref. Mor07 was used). Table 2 lists the reaction, absorption, and (elastic) breakup cross sections from the CDCC calculations. Clearly, inclusion of the continuum has a much larger effect on the elastic cross section for 6He. This is in apparent contraction with the fact that the elastic breakup cross section is larger in the 8B case, as shown in the Table. The apparent inconsistency can be explained by looking at the absorption and reaction cross sections. In the 8B reaction, the inclusion of the continuum couplings produces only a small increase of the absorption cross section, and the reaction cross section is essentially increased by the elastic breakup cross section. By contrast, in the 6He reaction, the absorption cross section increases by almost a factor of 3 when adding the continuum couplings and this results also in an equivalent increase of the reaction cross section, hence the enhanced effect on the elastic cross section. One may anticipate that the increased absorption cross section for 6He is associated with processes in which the valence neutrons experience some kind of nonelastic interaction with the target, such as transfer to bound states.
@@
Figure 17: Comparison of one-channel (i.e., no continuum) calculations and full CDCC calculations for (a) 8B+90Zr and (b) 6He+208Pb at near-barrier energies. See text for details.
@@
3.7.1 Two-body observables
The solution of the coupled equations provides the $S-$matrix elements connecting the ground state with the continuum states, from which breakup observables can be constructed. In principle, these breakup $S-$matrix elements corresponding to a given set of quantum numbers would be a continuous function of the excitation energy (or momentum). However, due to the discretization procedure, the S-matrices calculated in CDCC correspond to discrete values of the energy. In the binning method, an approximation to this continuous $S-$matrix can be obtained dividing the discrete $S-$matrix by the square root of the bin width, i.e.,
@@ 3.7.1 Two-body observables
3.7.2 Three-body observables
In XCDCC, the solution of the coupled-channel equations provides some discrete breakup transition amplitudes, denoted $T_{M_{0},M^{\prime}}^{i,J_{0},J^{\prime}}(\theta_{i},K_{i})$, connecting an initial state $|J_{0}M_{0}\rangle$ with a three-body final state comprised by the target (assumed to be structureless), the valence particle and the core, and some discrete value of the final projectile-target c.m. momentum $\vec{K}_{i}=\{\theta_{i},K_{i}\}$. The first step of the formalism is to relate these discrete amplitudes with the actual breakup scattering amplitudes, without continuum discretization, that we denote as $T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})$, where $\vec{k}$ is the projectile internal relative momentum. Formally, these breakup transition amplitudes can be written in integral form as:
@@ 3.7.2 Three-body observables
In order to relate the discrete and continuous amplitudes, one can approximate the exact wavefunction $\Psi_{J_{0},M_{0}}$ in the equation above by its XCDCC counterpart and introduce the approximate completeness relation in the truncated discrete basis $\{\phi^{(N)}_{n,J,M};i=1,\ldots,N\}$:
@@
The three-body observables, assuming the energy of the core is measured, are given by Tos01 :
@@
An application of this formalism is presented in Fig. 19, corresponding to the breakup of 11Be on a proton target at $E_{p}=63.7$ MeV/u. The left panel corresponds to the differential cross section with respect to the final $n$+10Be relative energy and the right panel to the breakup with respect to the final 10Be energy (integrated in the 10Be and neutron angles). These observables were computed by transforming the XCDCC breakup transition amplitudes by means of Eqs. (36) and (37), and integrating over the unmeasured variables. The XCDCC calculations were performed with a 11Be model including the 10Be ground ($0^{+}$) and first excited ($2^{+}$) states. The formalism allows to separate and quantify the contribution of these two states of 10Be. Note that, due to energy conservation, in the relative energy distribution the contribution coming from the 10Be($2^{+}$) state contributes only above the excitation energy of this state ($E_{x}=3.367$ MeV).
Figure 19: Differential breakup cross sections, with respect to the $n$-10Be relative energy (a) and with respect to the final 10Be energy (b), for the breakup of 11Be on protons at 63.7 MeV/nucleon. The contributions of the 10Be ground ($0^{+}$) and first excited ($2^{+}$) states are shown in each panel. In the left panel, the result of the calculation neglecting the 10Be excitation mechanism (labelled “no DCE”) is also shown. The vertical arrow denotes the threshold for 10Be($2^{+}_{1}$)+n. Figure adapted from Ref. Die17 .
@@ 3.7.3 Pseudostates versus bins
The PS discretization method becomes particularly suitable when dealing with narrow resonances. In the case of bins, one needs fine discretization in order to obtain a detailed description of the resonance region in the relative energy spectrum. By contrast, using PS’s one can obtain a detailed description of the resonance profile with a relatively small basis, with the aid of the convolution procedure discussed in Sec. 3.7.2. This is illustrated in Fig. 20, where we show the relative energy differential cross section corresponding to the 11Be+p $\to$ 10Be+n+p reaction at an incident energy of 63.7 MeV/nucleon. The 11Be bound and continuum states were generated with the structure model of Ref. Cap04 , which does not account for 10Be excitations. Continuum waves $s_{1/2}$, $p_{1/2}$, $p_{3/2}$, $d_{3/2}$, $d_{5/2}$ were included. For the CDCC-Bin calculation, the continuum was truncated at $E_{\mathrm{rel}}=12$  MeV and divided into 12 bins for each partial wave, evenly spaced in momentum. The relative-energy differential cross section was obtained by dividing the cross section of each bin by the width of that bin [see Eq. (31)]. The CDCC-PS calculations employ a THO basis with $N=30$ ($N=25$) oscillator functions for $\ell=1,2$ ($\ell=0$). After diagonalization of the 11Be Hamiltonian on this basis, only the eigenvalues below 12 MeV were retained. This results in 13 states for each partial wave (15 for $\ell=0$). The calculated breakup scattering amplitudes were then convoluted using Eq. (35), from which the differential cross sections, Eq. (36) was then evaluated. Finally, the latter was integrated over the solid angles $\Omega_{k}$ and $\Omega_{K}$.
@@
However, there are other situations where the binning procedure turns out to be more efficient than the PS method. One such situation is the case of reactions of highly polarizable weakly-bound nuclei (such as neutron-halo nuclei) on high-Z targets. These reactions are characterized by the presence of strong long-range Coulomb couplings, which tend to emphasize low-lying excitation states and probe large separations. In this case, the binning method has been shown to provide more stable results with respect to the basis size.

### v2

@@ 3.7 Microscopic CDCC
The standard CDCC method assumes a cluster (two-body or three-body) description of the projectile nucleus. This simplification has of course limitations and drawbacks. For example: (i) it requires cluster-target optical potentials, which are not always well determined; (ii) the extension to more than three bodies is very challenging and currently not available; (iii) excitations of the fragments are ignored altogether or, at most, approximately included with some collective model. To overcome these problems, a microscopic version of the CDCC method (MCDCC) has been proposed by Descouvemont and co-workers Des13 ; Des18 . The method uses a many-body description of the projectile states, based on a cluster approximation, known as resonating group method (RGM). In the RGM, an eigenstate of the projectile Hamiltonian is written as an antisymmetric product of cluster wave functions. For example, for a 7Li projectile, described as $\alpha+t$, the RGM wave function is expressed as:
@@
where $\phi_{\alpha}$ and $\phi_{t}$ are shell model wave functions of the $\alpha$ and $t$ clusters, $\ell$ their relative orbital angular momentum and $J$ the total spin. In Eq. (33), $\vec{\rho}$ is the relative coordinate (see Fig. 19), and ${\mathcal{A}}$ is the 7-body antisymmetrization operator which takes into account the Pauli principle among the 7 nucleons of the projectile. The function $g^{\ell J}_{i}(\rho)$ is determined from a Schrödinger equation associated with the projectile Hamiltonian. Continuum states are included using a pseudo-state basis.
@@
Figure 19 shows an application of the method to the reaction 7Li+208Pb at near-barrier energies. Experimental data are compared with a one-channel calculation (only 7 Li g.s.), two-channel calculation (ground plus first excited state) and several CDCC calculations including continuum states up to a certain projectile angular momentum. It is seen that a good description of the data is achieved when a sufficiently large number of continuum states are included.
@@
Other applications of the microscopic CDCC method include its extension for three-body projectiles including a single nucleon as one of the three particles, such as 8Li or 8B as $(^{4}\text{He}+t+n)$ and $(^{4}\text{He}+{{}^{3}}\text{He}+p)$, respectively Descouvemont2018 . Figure 20 shows results for the elastic scattering of 7Li on 12C and 209Bi targets at low energies. The dashed line corresponds to the calculation including only the projectile ground state, while the solid line is the full microscopic CDCC calculations that exhibit a better agreement with the data.
@@ 3.8 Interpretation of the CDCC wavefunction and cross sections
For a meaningful comparison with experimental data, it is important to properly understand which information and observables are (and which are not!) provided by the CDCC wavefunction. Of course, elastic scattering is the most direct outcome of the method and the one that can be compared with data more easily. By construction, the elastic scattering produced by the CDCC calculation incorporates the effect of the coupling to the inelastic and breakup channels included in the CDCC modelspace. According to our former notation, these breakup channels correspond to the three-body final states of the form $b+x+A$ in which the target and the projectile subsystems $b$ and $x$ escape the interaction region and remain in their ground state. Morover, in addition to these channels explicitly included in the modelspace, the CDCC includes effectively the effects of other channels associated with possible excitations of the target and/or of the fragments via the optical potentials.
@@
When the binning method is employed, an approximate double differential cross section (with respect to the c.m. angle and relative energy) can be obtained by dividing each angular distribution by the bin width, as follows:
@@
Regarding the nonelastic breakup contributions, which are effectively included in the fragment-target imaginary potentials, the CDCC method by itself does not provide detailed breakup cross sections involving excitations. However, one can define and compute a total absorption cross section that accounts for their overall contribution. For that, one first derives a reaction cross section in terms of the CDCC S-matrices as [^1] See Ref. Des17 for a different convention.:
$\sigma_{\text{reac}}$ $=\frac{\pi}{K^{2}_{i}}\sum_{J_{T}\pi}\sum_{L_{i}}\frac{2J_{T}+1}{(2J^{i}_{p}+1)(2J^{i}_{t}+1)}\left(1-|S^{J_{T}}_{\beta_{i},\beta_{i}}|^{2}\right),$ (37)
@@
$\sigma_{\text{bu}}$ $=\frac{\pi}{K^{2}_{i}}\frac{K}{K_{i}}\sum_{J_{T}\pi}\sum_{LL_{i}}\frac{2J_{T}+1}{(2J^{i}_{p}+1)(2J^{i}_{t}+1)}|S^{J_{T}}_{\beta,\beta_{i}}|^{2}.$ (38)
Then, one may define an absorption cross section as $\sigma_{\text{abs}}=\sigma_{\text{reac}}-\sigma_{\text{bu}}$. This absorption cross section will account for all nonelastic processes associated with the imaginary part of the fragment-target potentials, such as target excitation (i.e., inelastic scattering), complete fusion (i.e., capture of the whole projectile) or incomplete fusion (i.e., capture of one of the projectile fragments). Very often, comparison with experimental data requires more detailed cross sections, such as the angular or energy distributions of the fragment that survives after the capture (absorption) of the other fragment. As noted above, these observables are not directly provided by the CDCC method, at least in its standard formulation. However, as will be shown in Sec. 3.8.2, the CDCC wavefunction can be used as an input for a suitable formalism capable of providing such observables.
The absorption cross section plays an important role in understanding the CDCC results. To quantify the importance of the continuum on the elastic cross sections it is common to compare full CDCC results with no-continuum calculations, in which only the ground-state to ground-state coupling potential is retained. An example is shown in Fig. 21, which depicts the results for the 6He+208Pb reaction at 18 MeV and the 8B+90Zr reaction at 26.5 MeV. Dashed and solid lines denote the no-continuum and full CDCC calculations, respectively (in the 6He case, the improved dineutron model of Ref. Mor07 was used). Table 2 lists the reaction, absorption, and (elastic) breakup cross sections from the CDCC calculations. Clearly, inclusion of the continuum has a much larger effect on the elastic cross section for 6He. This is in apparent contradiction with the fact that the elastic breakup cross section is larger in the 8B case, as shown in the Table. The apparent inconsistency can be explained by looking at the absorption and reaction cross sections. In the 8B reaction, the inclusion of the continuum couplings produces only a small increase of the absorption cross section, and the reaction cross section is essentially increased by the elastic breakup cross section. By contrast, in the 6He reaction, the absorption cross section increases by almost a factor of 3 when adding the continuum couplings and this results also in an equivalent increase of the reaction cross section, hence the enhanced effect on the elastic cross section. One may anticipate that the increased absorption cross section for 6He is associated with processes in which the valence neutrons experience some kind of nonelastic interaction with the target, such as transfer to bound states. These processes will be discussed in a subsequent section.
@@
Figure 21: Comparison of one-channel (i.e., no continuum) calculations and full CDCC calculations for the elastic scattering of (a) 8B+90Zr and (b) 6He+208Pb at near-barrier energies. See text for details.
@@
3.8.1 Smoothing procedure for two-body observables
Owing to the discretization procedure inherent to the CDCC method, the breakup cross sections are given only for the discrete energies of the bins or pseudostates. In reality, breakup cross sections should be continuous functions of the relative energy between the fragments. Several procedures have been proposed to convert the discrete breakup energy distributions into continuous distributions for arbitrary values of the inter-fragment relative energy. We discuss here two such methods, one based on the S-matrices and the other on the scattering amplitudes.
The S-matrices calculated in CDCC correspond to discrete values of the energy. In the binning method, an approximation to this continuous $S-$matrix can be obtained dividing the discrete $S-$matrix by the square root of the bin width, i.e.,
@@ 3.8.1 Smoothing procedure for two-body observables
From the S-matrices, the double differential cross section for a $b-x$ breakup state with orbital angular momentum $\ell$ and wavenumber $k$ is given by:
$\frac{d^{2}\sigma_{\ell}(k)}{dk\,d\Omega_{\text{c.m.}}}=\sum_{m=0}^{\ell}\frac{\pi}{K_{0}^{2}}\left|\sum_{J,L}(2J+1)\langle\ell mL-m|J0\rangle|Y_{L-m}(\Omega_{\alpha})S^{J_{T}}_{\beta,\beta_{i}}(k)\right|^{2},$ (41)
@@
The connection between the discrete and continuous breakup cross sections can be also done at the level of the transition amplitudes themselves. In both CDCC and XCDCC, the solution of the coupled-channel equations provides some discrete breakup transition amplitudes, denoted $T_{M_{0},M^{\prime}}^{i,J_{0},J^{\prime}}(\theta_{i},K_{i})$, connecting an initial state $|J_{0}M_{0}\rangle$ with a three-body final state comprised of the target (assumed to be structureless), the valence particle and the core, at some discrete value of the final projectile-target c.m. momentum $\vec{K}_{i}=\{\theta_{i},K_{i}\}$. The first step of the formalism is to relate these discrete amplitudes with the actual breakup scattering amplitudes, without continuum discretization, that we denote as $T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})$, where $\vec{k}$ is the projectile internal relative momentum. Formally, these breakup transition amplitudes can be written in integral form as:
@@
In order to relate the discrete and continuous amplitudes, one can approximate the exact wavefunction $\Psi_{J_{0},M_{0}}$ in the equation above by its (X)CDCC counterpart and introduce the approximate completeness relation in the truncated discrete basis $\{\phi^{(N)}_{n,J,M};i=1,\ldots,N\}$:
@@
3.8.2 Three-body observables
Using the two-body transition amplitudes discussed in the previous subsection, the three-body observables, assuming the energy of the core is measured, are given by Tos01 :
@@ 3.8.2 Three-body observables
An application of this formalism is presented in Fig. 23, corresponding to the breakup of 11Be on a proton target at $E_{p}=63.7$ MeV/u. The top panel corresponds to the differential cross section with respect to the final $n$+10Be relative energy and the bottom panel to the breakup with respect to the final 10Be energy (integrated in the 10Be and neutron angles). These observables were computed by transforming the XCDCC breakup transition amplitudes by means of Eqs. (43)-(45), and integrating over the unmeasured variables. The XCDCC calculations were performed with a 11Be model including the 10Be ground ($0^{+}$) and first excited ($2^{+}$) states. The formalism allows to separate and quantify the contribution of these two states of 10Be. Note that, due to energy conservation, in the relative energy distribution the contribution coming from the 10Be($2^{+}$) state contributes only above the excitation energy of this state ($E_{x}=3.367$ MeV). The importance of the 10Be excitation during the reaction due to its interaction with the target nucleus is illustrated also in the top panel by means of a calculation in which the deformed part of the 10Be+p interaction is omitted (dot-dashed line).
Figure 23: Differential breakup cross sections, with respect to the $n$-10Be relative energy (a) and with respect to the final 10Be energy (b), for the breakup of 11Be on protons at 63.7 MeV/nucleon. The contributions of the 10Be ground ($0^{+}$) and first excited ($2^{+}$) states are shown in each panel. In the top panel, the result of the calculation neglecting the 10Be excitation mechanism (labelled “no DCE”) is also shown. The vertical arrow denotes the threshold for 10Be($2^{+}_{1}$)+n. Figure adapted from Ref. Die17 .
@@ 3.8.3 Pseudostates versus bins
The PS discretization method becomes particularly suitable when dealing with narrow resonances. In the case of bins, one needs fine discretization in order to obtain a detailed description of the resonance region in the relative energy spectrum. By contrast, using PS’s one can obtain a detailed description of the resonance profile with a relatively small basis, with the aid of the convolution procedure discussed in Sec. 3.8.2. This is illustrated in Fig. 24, where we show the 10Be+n relative energy differential cross section corresponding to the 11Be+p $\to$ 10Be+n+p reaction at an incident energy of 63.7 MeV/nucleon. The 11Be bound and continuum states were generated with the structure model of Ref. Cap04 , which does not account for 10Be excitations. Continuum waves $s_{1/2}$, $p_{1/2}$, $p_{3/2}$, $d_{3/2}$, $d_{5/2}$ were included. For the CDCC-Bin calculation, the continuum was truncated at $E_{\mathrm{rel}}=12$  MeV and divided into 12 bins for each partial wave, evenly spaced in momentum. The relative-energy differential cross section was obtained by dividing the cross section of each bin by the width of that bin [see Eq. (36)]. The CDCC-PS calculations employ a THO basis with $N=30$ ($N=25$) oscillator functions for $\ell=1,2$ ($\ell=0$). After diagonalization of the 11Be Hamiltonian on this basis, only the eigenvalues below 12 MeV were retained. This results in 13 states for each partial wave (15 for $\ell=0$). The calculated breakup scattering amplitudes were then convoluted using Eq. (43), from which the differential cross sections, Eq. (44), were then evaluated. Finally, the latter were integrated over the solid angles $\Omega_{k}$ and $\Omega_{K}$.
@@
However, there are other situations where the binning procedure turns out to be more efficient than the PS method. One such situation is the case of reactions of highly polarizable weakly-bound nuclei (such as neutron-halo nuclei) on high-Z targets. These reactions are characterized by the presence of strong long-range Coulomb couplings, which tend to emphasize low-lying excitation states and probe large separations. In this case, the binning method has been shown to provide more stable results with respect to the basis size Rod11 .